IN HOW MANY WAYS CAN 6 INDIVIDUALS BE SEATED IN A ROUND TABLE WITH 6 CHAIRS?
A.) SUPPOSE 2 PERSONS WANTED TO BE SEATED SIDE BY SIDE ,IN HOW MANY WAYS CAN THEY DO IT?
B.) IN HOW MANY WAYS CAN THESE 6 INDIVIDUALS ARRAGE THEMSELVES IF 2 AMONG THEM REFUSE TO SIT TOGETHER?

Answer by ikleyn(53751)

.
1)  IN HOW MANY WAYS CAN 6 INDIVIDUALS BE SEATED IN A ROUND TABLE WITH 6 CHAIRS?


    In  5! = 1*2*3*4*5 = 120 ways.    (The "ways" are circular permutations in this case (!) )




2)  SUPPOSE 2 PERSONS WANTED TO BE SEATED SIDE BY SIDE ,IN HOW MANY WAYS CAN THEY DO IT ?

    
    In 2*4! = 2*(1*2*3*4) = 2* 24 = 48 ways.


    The factor "2" before 4! is due to the fact that there are 2 permutations of these two persons.




3)  IN HOW MANY WAYS CAN THESE 6 INDIVIDUALS ARRAGE THEMSELVES IF 2 AMONG THEM REFUSE TO SIT TOGETHER?


    This number is the complement of 48 to 120:  in 120-48 = 72 ways.

-------------------

See the lesson
    - Persons sitting around a cicular table
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.